\int \:\left(x^2-2\right)^3x^3dx
1
Expert's answer
2020-10-01T16:01:07-0400

"\int(x^2 - 2)^3x^3dx = \begin{bmatrix}
u = x^2 - 2 \\
du = 2xdx \\
dx = \frac{1}{2x}du
\end{bmatrix}
="

"= \frac{1}{2} \int u^3(u+2)du = \frac{1}{2} \int (u^4 + 2u^3)du = \\ = \frac{1}{2}(\int u^4du + 2\int u^3 du) = \frac{1}{2}(\frac{u^5}{5} + \frac{u^4}{2}) = \\ =\frac{u^5}{10} + \frac{u^4}{4} = \begin{bmatrix}
u = x^2 - 2
\end{bmatrix} 
= \frac{(x^2 - 2)^5}{10} + \frac{(x^2 - 2)^4}{4} + C"